Which of this is not true?

a.\\(i \times i=j \times j=k \times k=0\\)

b.\\(\alpha \times \beta = \beta \times \alpha \\)

c.\\(\alpha \times \beta =- \beta \times \alpha \\)

d.\\(\alpha \times \beta \neq \beta \times \alpha \\)


1
Expert's answer
2020-03-30T12:34:23-0400

As per the given question

a. "\hat{i} \times \hat{i}=\hat{j} \times \hat{j}=\hat{k} \times \hat{k}=0"

we know that "\hat{i}\times\hat{i}=\sin\theta\hat{n}"

Here angle between i and i, j and j , k and k are 0, hence option A is correct.

b) "\alpha \times \beta = \beta \times \alpha"

we know that

"\alpha\times \beta" "=-\beta\times \alpha" Hence option b is false.

c) "-\beta\times \alpha"

From the above we can say that option c is correct.

d)"\alpha \times \beta \neq \beta \times \alpha"

from the option above, we can say that option d is also correct.


Hence option b is incorrect